OpenStudy (anonymous):

whats 1.17 repeating as a fraction

OpenStudy (joannablackwelder):

First, write as an equation x=1.17 repeating

OpenStudy (joannablackwelder):

Multiply by a power of 10 on both sides to move the repeating digits to the left of the decimal place. 100x = 117.17 repeating

OpenStudy (joannablackwelder):

Now, subtract the first equation from the second. 100x = 117.17 - x = 1.17

OpenStudy (joannablackwelder):

99x=116

OpenStudy (joannablackwelder):

x=116/99

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

1.171717..... 1+X=1.1717... So X=0.1717... Let X*100=17.1717... And 100X-17=0.1717...=X 99X=17 X=17/99=0.1717... So 1.1717...=1+17/99= 116/99